OpenStudy (anonymous):

prove: [(3x+2)/x+1]-5/6=0

OpenStudy (anonymous):

[(3x+2)/x+1]=5/6 3x+2=5x+5/6 6/6(3x+2)=(5x+5)/6 (18x+12)/6=(5x+5)/6 (13x+7)/6=0 13x=-7 x=-7/13 [[3(-7/13)+2]/[(-7/13)+1]-(5/6)=0

OpenStudy (anonymous):

I dropped that negative sign in the next to last step... Needless to say that problem took a lot longer than it should have...

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

oh my i've typed it wrong. it's supposed to be x-1... :( [(3x+2)/x-1]-5/6=0

OpenStudy (anonymous):

how is it now? please help me..

OpenStudy (anonymous):

[(3x+2)/x-1]=5/6 3x+2=5x-5/6 6/6(3x+2)=(5x-5)/6 (18x+12)/6=(5x-5)/6 (13x+17)/6=0 13x=-17 x=-17/13 [[3(-17/13)+2]/[(-17/13)+1]-(5/6)=0